# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v3

@@ Abstract.
We study abelian surfaces defined over finite fields which do not contain any possibly singular curve of genus less than or equal to $3$. Firstly, we complete and expand the characterisation of isogeny classes of abelian surfaces with no curves of genus up to $2$ initiated by the first author et al. in previous work. Secondly, we show that, for simple abelian surfaces, containing a curve of genus $3$ is equivalent to admitting a polarisation of degree $4$. Thanks to this result, we can use existing algorithms to check which isomorphism classes in the isogeny classes containing no genus $2$ curves have a polarisation of degree $4$. Thirdly, we characterise isogeny classes of abelian surfaces with no curves of genus $\leq 2$, containing no abelian surface with a polarisation of degree $4$. Finally, we describe the absolutely irreducible genus $3$ curves lying on abelian surfaces containing no curves of genus less than or equal to $2$, and show that their number of rational points is far from the Serre–Weil bound.
@@ Introduction
From now on, we direct our attention to the case of finite fields. Isogeny classes in which no abelian surface admits a principal polarisation are classified in [15]. Furthermore, even if an isogeny class is principally polarisable, it might not contain a Jacobian surface nor the product of two elliptic curves [16]. Indeed, by a theorem of Weil [16, Thm. 1.3], we have one more option among principally polarisable abelian surfaces, that is, Weil restrictions of elliptic curves defined over a quadratic extension of the finite field. To sum up, studying the minimal genus of algebraic curves lying over an abelian surface defined over a finite field is a hard question, and it is in fact open in its full generality.
@@
    $A$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$.
@@
Secondly, we consider the irreducible Weil polynomials $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. When the corresponding isogeny class $\mathcal{C}$ is ordinary, the third author designed an algorithm in [19] to compute the isomorphism classes of abelian surface in $\mathcal{C}$ admitting a polarisation of degree $4$, and hence containing an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$. See Section 6 for an overview and examples. By applying Main Theorem 2 to our isogeny classes, we easily obtain the following statements about curves of arithmetic genus $\leq 3$ (and not just $=3$). See Corollary 2.5 for the proof.
@@
($\bigstar$) There is no $\text{There is no }A\text{ in }\mathcal{C}\text{ admitting a polarisation of degree 4.}$.
@@
    Assume $\mathcal{C}$ is ordinary. Then (italic-$\bigstar$3) holds if and only if there is no $A$ in $\mathcal{C}$ with maximal $\mathbb{F}_{q}$-endomorphism ring admitting a polarisation of degree $4$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then (italic-$\bigstar$3) holds if and only if $2$ is inert in $K^{+}$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, so that $f(t)=t^{4}+bt^{2}+q^{2}$. Then (italic-$\bigstar$3) is equivalent to
    • $b=1-2q$ and $q$ is odd, if $\mathcal{C}$ is ordinary,
@@
In Section 1, we prove Main Theorem 1, expanding on [2] and characterising the isogeny classes of abelian surfaces containing no absolutely irreducible curves of geometric genus smaller than or equal to $2$. In Section 2, we start the characterisation of abelian surfaces containing no curves of genus $3$ among those containing no curves of genus up to $2$. In particular, in Theorem 2.2, we prove the key equivalence between containing a curve of arithmetic genus $3$ and having a polarisation of degree $4$ stated above as Main Theorem 2. We derive also some consequences stated above in the Corollary. In Section 3 we collect technical results on the factorisation of $2$ in the extension $K/\mathbb{Q}$, that we shall use throughout the paper. Section 4 is devoted to results on kernels of polarisations of abelian varieties, based on the work of Howe [13, 14]. In Section 5, we prove an expanded version of Main Theorem 3, namely Theorem 5.7, in which we give necessary and sufficient conditions on an isogeny class to not contain surfaces admitting a polarisation of degree $4$. Section 6 describes the algorithm proposed by the third author in [19] to compute the isomorphism classes of abelian surface in an isogeny class admitting a polarisation of degree $4$, and provides examples for our Theorem 5.7. Finally, in Section 7, we characterise absolutely irreducible smooth curves of genus $3$ lying on abelian surfaces containing no curves of genus less than $2$, and give bounds for their number of rational points, showing that such genus $3$ curves are far from being maximal.
@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
By a curve we mean a possibly singular one-dimensional projective variety defined over $\mathbb{F}_{q}$. We will only consider $\mathbb{F}_{q}$-irreducible curves in this paper. In some cases, we will need our curves to be absolutely irreducible, that is, irreducible over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$. We say that a curve $C$ defined over $\mathbb{F}_{q}$ lies on an abelian surface $A$ or that $A$ contains $C$, if $C$ is a closed subvariety of $A$, that is, there is an $\mathbb{F}_{q}$-embedding of $C$ in $A$. A divisor of $A$ defined over $\mathbb{F}_{q}$, or a rational divisor over $\mathbb{F}_{q}$, is a formal sum, with integer coefficients, of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ and lying on $A$ [^1] A divisor is often called rational over $\mathbb{F}_{q}$ when it is invariant under the action of $\mathrm{Gal}(\overline{\mathbb{F}}_{q}/\mathbb{F}_{q})$. This implies in particular that the divisor can be written as a formal sum of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ (see [9, §3.1]). In the present paper we shall only use this characterization, hence we take it as a definition.. It is called effective if all its coefficients are non-negative and ample if a multiple of it is very ample, that is, defines an embedding of the surface in a projective space.
The arithmetic genus $p_{a}(C)$ of an $\mathbb{F}_{q}$-irreducible curve $C$ is defined as $1-\chi(C)$, $\chi(C)$ being the Euler-Poincaré characteristic of $C$ [25, Ch. IV, Sec. 2]. Consider the absolutely irreducible components $C_{i}$ of $C$, and their normalisations $\tilde{C_{i}}$. We define the geometric genus $g(C)$ of $C$ as $g(C)=p_{a}(\tilde{C_{1}})+\ldots+p_{a}(\tilde{C_{n}})$. We have that $g(C)\leq p_{a}(C)$ with an equality if and only if $C$ is smooth. If $C$ is an $\mathbb{F}_{q}$-irreducible curve lying on an abelian surface $A$ then, since the class of the canonical divisor of $A$ is trivial, we have that $C^{2}=2p_{a}(C)-2$ by the adjunction formula [25, Ch. IV, Sec. 2, Prop. 5].
The goal of this section is to prove some characterisations and properties of Weil polynomials of abelian surfaces which do not contain absolutely irreducible curves of geometric genus 00, $1$ or $2$. We start with the following useful lemma.
@@
[…]
@@
    $A$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $2$.
@@
    $A$ does not contain absolutely irreducible curves of geometric genus $0,1$ or $2$;
@@
    $A$ is simple and not isogenous to the Jacobian of an absolutely irreducible smooth genus $2$ curve;
@@
      $A$ is not isogenous to a principally polarised abelian surface, which is equivalent to have $a^{2}-b=q$, $b<0$ and all prime divisors of $b$ are congruent to $1\bmod 3$;
@@
      $A$ is isogenous to a Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ (hence $a=0$) and one of the following conditions holds:
      • $b=1-2q$;
      • $p>2$ and $b=2-2q$;
      • $p\equiv 11\bmod 12$, $q$ is a square and $b=-q$;
      • $p=3$, $q$ is a square and $b=-q$;
      • $p=2$, $q$ is nonsquare and $b=-q$;
      • $q=2$ and $b=-4$;
      • $q=3$ and $b=-6$.
@@
We start by proving (i)$\Rightarrow$(ii). By assumption, $A$ does not contain an elliptic curve, hence it is simple. Assume that $A$ is isogenous to the Jacobian of a smooth absolutely irreducible genus $2$ curve $C$. Since $C$ is canonically embedded into its Jacobian, Lemma 1.1 would imply that $A$ contains an absolutely irreducible curve of geometric genus $\leq 2$, in contradiction with (i).
Now we show that (ii)$\Rightarrow$(i). First note that $A$ cannot contain an absolutely irreducible curve $D$ of geometric genus 00. Indeed, the normalisation $\tilde{D}$ of $D$ is birationally equivalent to $\mathbb{P}^{1}$ and, by [20, Cor. 3.8], the only rational maps from $\mathbb{P}^{1}$ to $A$ are the constant maps. The abelian variety $A$ cannot contain an absolutely irreducible curve of geometric genus $1$ as well. Indeed, the normalisation of such a curve would be an elliptic curve (see for example [9, Prop. 1]) and hence an isogeny factor of $A$. Similarly, $A$ cannot contain an absolutely irreducible curve $D$ of geometric genus $2$ because the induced map $\tilde{D}\to A$ would give a morphism $\mathop{\rm Jac}\nolimits(\tilde{D})\to A$, which is an isogeny since $A$ is simple.
We now focus on (ii)$\Rightarrow$(iii). We distinguish two cases. Assume first that $A$ is not isogenous to a principally polarised abelian surface. This corresponds to (iii)(iii).(a). The equivalence in terms of the coefficients of the Weil polynomial follows from [15, Thm. 1]. Assume now that $A$ is isogenous to a principally polarised abelian surface. By a classification theorem due to Weil (see for instance [16, Thm. 1.3]), a principally polarised abelian surface defined over $\mathbb{F}_{q}$ is exactly one of the following: a product of two elliptic curves defined over $\mathbb{F}_{q}$ with the product polarisation; the Jacobian of an absolutely irreducible genus $2$ curve defined over $\mathbb{F}_{q}$ with the canonical polarisation; the Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ with the induced polarisation. Since we are assuming (ii), we exclude the first two cases. So, $A$ is the Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$. Finally, the characterisation of the coefficients of the Weil polynomial follows from [2, Prop. 4.3] and the beginning of its proof. Hence, we are in case (iii)(iii).(b).
We conclude by showing that (iii)$\Rightarrow$(ii). For $A$ satisfying (iii)(iii).(a) the implication is clear. If $A$ satisfies (iii)(iii).(b) the result follows from [16, Thm. 1.2-(2), Table 1.2]. ∎
@@
    $A$ is simple and not isogenous to a Jacobian surface, that is, the Jacobian of an absolutely irreducible smooth genus $2$ curve;
@@
    $A$ does not contain absolutely irreducible curves of arithmetic genus 00, $1$ or $2$;
@@
    $A$ does not contain absolutely irreducible smooth curves of genus 00, $1$ or $2$.
@@
Note that one can distinguish isogeny classes of abelian surfaces admitting no principal polarisation and simple isogeny classes of Weil restrictions by their splitting behavior. From [15, p. 122], we know that an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is in a not principally polarisable isogeny class splits over the cubic extension of the base field, that is $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{3}}$ is not simple. If $A$ is a simple abelian surface over $\mathbb{F}_{q}$ then $A$ is isogenous to a Weil restriction of an elliptic curve defined over $\mathbb{F}_{q^{2}}$ if and only if $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is not simple; see [15, Lemma 4].
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$ in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Then $A$ is either ordinary or supersingular. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ then $A$ is ordinary if and only if $a\neq 0$. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, then $A$ is ordinary if and only if $b=1-2q$ or, $b=2-2q$ and $p>2$.
@@
Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Note that $A$ cannot have $p$-rank $1$ since in that case it would be principally polarisable by [18, Thm. 4.3] [^2] The authors of [18] call the neither ordinary nor supersingular case “mixed”.. If $a=0$ then we have $b=a^{2}-q=-q$, and therefore $A$ is supersingular. Suppose $a\neq 0$. By [15, Thm. 2], the abelian variety $A$ cannot be supersingular. Hence, $A$ is ordinary. Now, assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Obviously if $b=1-2q$, or $b=2-2q$ and $p>2$ then $A$ is ordinary. By Theorem 1.4.(iii)(iii).(b), those are the only possible values for $b$ to have ordinarity. To conclude the proof, we need to show that if $A$ is non-ordinary then it is supersingular. Say that $A$ is the Weil restriction of an elliptic curve $E$ over $\mathbb{F}_{q^{2}}$. If $A$ is non-ordinary, then $E$ is supersingular, and hence $A$ is also supersingular. ∎
@@
Let $\alpha,q/\alpha,\beta,q/\beta$ be the complex roots of $f(t)$. Then $-a=\alpha+q/\alpha+\beta+q/\beta$. If $a=0$ then either $\alpha=-\beta$ or $\alpha=-q/\beta$. In both cases $\beta\in\mathbb{Q}(\alpha)$, so the extension is normal and $K$ is Galois. If $a\neq 0$ then $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and $A$ is ordinary by Lemma 1.12. Hence $K$ is Galois by [13, Lemma 12.1]. ∎

### v4

@@ Abstract.
We study abelian surfaces defined over finite fields which do not contain any possibly singular curve of genus less than or equal to $3$. Firstly, we complete and expand the characterisation of isogeny classes of abelian surfaces with no curves of genus up to $2$ initiated by the first author et al. in previous work. Secondly, we show that, for simple abelian surfaces, containing a curve of genus $3$ is equivalent to admitting a polarisation of degree $4$. Thanks to this result, we can use existing algorithms to check which isomorphism classes in the isogeny classes containing no genus $2$ curves have a polarisation of degree $4$. Thirdly, we characterise isogeny classes of abelian surfaces with no curves of genus $\leq 2$, containing no abelian surface with a polarisation of degree $4$. Finally, we describe the absolutely irreducible genus $3$ curves lying on abelian surfaces containing no curves of genus less than or equal to $2$.
@@ Introduction
From now on, we direct our attention to the case of finite fields. Isogeny classes in which no abelian surface admits a principal polarisation are classified in [12]. Furthermore, even if an isogeny class is principally polarisable, it might not contain a Jacobian surface nor the product of two elliptic curves [13]. Indeed, by a theorem of Weil [13, Thm. 1.3], we have one more option among principally polarisable abelian surfaces, that is, Weil restrictions of elliptic curves defined over a quadratic extension of the finite field. To sum up, studying the minimal genus of algebraic curves lying over an abelian surface defined over a finite field is a hard question, and it is in fact open in its full generality.
@@
    AA contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$.
@@
Secondly, we consider the irreducible Weil polynomials $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. When the corresponding isogeny class $\mathcal{C}$ is ordinary, the third author designed an algorithm in [19] to compute the isomorphism classes of abelian surface in $\mathcal{C}$ admitting a polarisation of degree $4$, and hence containing an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$. See Section 6 for an overview and examples. By applying Main Theorem 2 to our isogeny classes, we easily obtain the following statements about curves of arithmetic genus $\leq 3$ (and not just =3=3). See Corollary 2.5 for the proof.
@@
[…]
@@
    Assume $\mathcal{C}$ is ordinary. Then ($\bigstar$3) holds if and only if there is no $A$ in $\mathcal{C}$ with maximal $\mathbb{F}_{q}$-endomorphism ring admitting a polarisation of degree $4$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Then ($\bigstar$3) holds if and only if $2$ is inert in $K^{+}$.
@@
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, so that $f(t)=t^{4}+bt^{2}+q^{2}$. Then ($\bigstar$3) is equivalent to
    • b=12qb=1-2q and $q$ is odd, if $\mathcal{C}$ is ordinary,
@@
In Section 1, we prove Main Theorem 1, expanding on [2] and characterising the isogeny classes of abelian surfaces containing no absolutely irreducible curves of geometric genus smaller than or equal to $2$. In Section 2, we start the characterisation of abelian surfaces containing no curves of genus $3$ among those containing no curves of genus up to $2$. In particular, in Theorem 2.2, we prove the key equivalence between containing a curve of arithmetic genus $3$ and having a polarisation of degree $4$ stated above as Main Theorem 2. We derive also some consequences stated above in the Corollary. In Section 3 we collect technical results on the factorisation of $2$ in the extension $K/\mathbb{Q}$, that we shall use throughout the paper. Section 4 is devoted to results on kernels of polarisations of abelian varieties, based on the work of Howe [15, 16]. In Section 5, we prove an expanded version of Main Theorem 3, namely Theorem 5.7, in which we give necessary and sufficient conditions on an isogeny class to not contain surfaces admitting a polarisation of degree $4$. Section 6 describes the algorithm proposed by the third author in [19] to compute the isomorphism classes of abelian surface in an isogeny class admitting a polarisation of degree $4$, and provides examples for our Theorem 5.7. Finally, in Section 7, we describe absolutely irreducible smooth curves of genus $3$ lying on abelian surfaces containing no curves of genus less than $2$. In particular, we give bounds for their number of rational points, showing that such genus $3$ curves are far from being maximal.
@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
By a curve we mean a possibly singular one-dimensional projective variety defined over $\mathbb{F}_{q}$. We will only consider $\mathbb{F}_{q}$-irreducible curves in this paper. In some cases, we will need our curves to be absolutely irreducible, that is, irreducible over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$. We say that a curve $C$ defined over $\mathbb{F}_{q}$ lies on an abelian surface $A$ or that $A$ contains $C$, if $C$ is a closed subvariety of $A$, that is, there is an $\mathbb{F}_{q}$-embedding of $C$ in $A$. A divisor of $A$ defined over $\mathbb{F}_{q}$, or a rational divisor over $\mathbb{F}_{q}$, is a formal sum, with integer coefficients, of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ and lying on AA [^1] A divisor is often called rational over $\mathbb{F}_{q}$ when it is invariant under the action of $\mathrm{Gal}(\overline{\mathbb{F}}_{q}/\mathbb{F}_{q})$. This implies in particular that the divisor can be written as a formal sum of $\mathbb{F}_{q}$-irreducible curves defined over $\mathbb{F}_{q}$ (see [9, §3.1]). In the present paper we shall only use this characterization, hence we take it as a definition.. It is called effective if all its coefficients are non-negative and ample if a multiple of it is very ample, that is, defines an embedding of the surface in a projective space.
[…]
The goal of this section is to prove some characterisations and properties of Weil polynomials of abelian surfaces which do not contain absolutely irreducible curves of geometric genus 0, 11 or $2$. We start with the following useful lemma.
@@
Set $D=f(C)$ and denote by $f_{C}\colon C\to D$ the restriction of $f$ to $C$. Since $f$ is an isogeny, then $D$ is an $\mathbb{F}_{q}$-irreducible curve. By the universal property of the normalisation $\tilde{D}\to D$, the composition of the normalisation $\tilde{C}\to C$ with $f_{C}\colon C\to D$ factors as the composition of a morphism $\tilde{f}_{C}\colon\tilde{C}\to\tilde{D}$ with $\tilde{D}\to D$. The morphism $\tilde{f}_{C}$ factors as $\tilde{C}\overset{i}{\to}Y\overset{s}{\to}\tilde{D}$ with $Y$ a normal scheme, $i$ a purely inseparable morphism and $s$ a separable morphism. Then, $g(\tilde{C})=g(Y)$. Finally, the Riemann-Hirwitz Theorem applied to $s:Y\to\tilde{D}$ gives us $g(\tilde{D})\leq g(Y)$. Combining, we obtain $g(C)=g(\tilde{C})=g(Y)\geq g(\tilde{D})=g(D)$.
@@
    AA contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $2$.
@@
    AA does not contain absolutely irreducible curves of geometric genus $0,1$ or $2$;
@@
    AA is simple and not isogenous to the Jacobian of an absolutely irreducible smooth genus $2$ curve;
@@
      AA is not isogenous to a principally polarised abelian surface, which is equivalent to have $a^{2}-b=q$, $b<0$ and all prime divisors of $b$ are congruent to $1\bmod 3$;
@@
      AA is isogenous to a Weil restriction of an elliptic curve defined over the quadratic extension of $\mathbb{F}_{q}$ (hence $a=0$) and one of the following conditions holds:
      • b=12qb=1-2q;
      • $p>2$ and b=22qb=2-2q;
      • $p\equiv 11\bmod 12$, $q$ is a square and b=qb=-q;
      • $p=3$, $q$ is a square and b=qb=-q;
      • $p=2$, $q$ is nonsquare and b=qb=-q;
      • $q=2$ and b=4b=-4;
      • $q=3$ and b=6b=-6.
@@
[…]
@@
    AA is simple and not isogenous to a Jacobian surface, that is, the Jacobian of an absolutely irreducible smooth genus $2$ curve;
@@
    AA does not contain absolutely irreducible curves of arithmetic genus 0, 11 or $2$;
@@
    AA does not contain absolutely irreducible smooth curves of genus 0, 11 or $2$.
@@
Note that one can distinguish isogeny classes of abelian surfaces admitting no principal polarisation and simple isogeny classes of Weil restrictions by their splitting behavior. From [12, p. 122], we know that an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is in a not principally polarisable isogeny class splits over the cubic extension of the base field, that is $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{3}}$ is not simple. If $A$ is a simple abelian surface over $\mathbb{F}_{q}$ then $A$ is isogenous to a Weil restriction of an elliptic curve defined over $\mathbb{F}_{q^{2}}$ if and only if $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is not simple; see [12, Lemma 4].
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$ in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Then $A$ is either ordinary or supersingular. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ then $A$ is ordinary if and only if $a\neq 0$. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, then $A$ is ordinary if and only if b=12qb=1-2q or, b=22qb=2-2q and $p>2$.
@@
Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Note that $A$ cannot have $p$-rank 11 since in that case it would be principally polarisable by [18, Thm. 4.3] [^2] The authors of [18] call the neither ordinary nor supersingular case “mixed”.. If $a=0$ then we have $b=a^{2}-q=-q$, and therefore $A$ is supersingular. Suppose $a\neq 0$. By [12, Thm. 2], the abelian variety $A$ cannot be supersingular. Hence, $A$ is ordinary. Now, assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Obviously if b=12qb=1-2q, or b=22qb=2-2q and $p>2$ then $A$ is ordinary. By Theorem 1.4.(iii)(iii).(b), those are the only possible values for $b$ to have ordinarity. To conclude the proof, we need to show that if $A$ is non-ordinary then it is supersingular. Say that $A$ is the Weil restriction of an elliptic curve $E$ over $\mathbb{F}_{q^{2}}$. If $A$ is non-ordinary, then $E$ is supersingular, and hence $A$ is also supersingular. ∎
@@
Let $\alpha,q/\alpha,\beta,q/\beta$ be the complex roots of $f(t)$. Then $a=\alpha+q/\alpha+\beta+q/\beta$. If $a=0$ then either $\alpha=-\beta$ or $\alpha=-q/\beta$. In both cases $\beta\in\mathbb{Q}(\alpha)$, so the extension is normal and $K$ is Galois. If $a\neq 0$ then $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and $A$ is ordinary by Lemma 1.12. Hence $K$ is Galois by [15, Lemma 12.1]. ∎